OpenStudy (anonymous):

can anyone help me with this problem please... 4d + 2d – 4d + 4 = 3d

OpenStudy (anonymous):

d=4 get all the variables on one side 4d+2d-4d+4=3d 4d+2d+4=3d+4d 4d+4=3d+4d-2d 4=3d+4d-2d-4d 4=7d-2d-4d 4=5d-4d 4=d

OpenStudy (anonymous):

the answer is D???/

OpenStudy (anonymous):

or its 4 I meant??

OpenStudy (anonymous):

yea at the bottom is the answer the rest is showing you how to get the answer step by step hope it helps

OpenStudy (anonymous):

all I needed was the step by step but thank you :)

OpenStudy (anonymous):

no problem!!